OpenStudy (anonymous):

(cos Θ − cos Θ)^2 + (cos Θ + cos Θ)^2

OpenStudy (anonymous):

i am pretty sure it is sin^2

OpenStudy (anonymous):

bu ti just need another opinion

OpenStudy (anonymous):

?

OpenStudy (lynfran):

(cos theta - cos theta)^2= 0^2 (cos theta + cos theta )^2 = (2cos theta)^2

OpenStudy (anonymous):

no, you just put that one..... and i said "?"

OpenStudy (mertsj):

\[(\cos \theta+\cos \theta)^2=\cos ^2\theta+2\cos ^2\theta+\cos ^2\theta=4\cos ^2 \theta\]

OpenStudy (anonymous):

so what type of cos equations would equal sin?

OpenStudy (anonymous):

and he is gone... ignore that last question...

OpenStudy (anonymous):

wish i i knew how to give medals...

OpenStudy (lynfran):

yes (2cos theta)^2 = 4 cos ^2 theta

OpenStudy (lynfran):

we have the Pythagorean identity cos^2 + sin ^2 =1 where cos^2 theta =1- sin^ theta